OpenStudy (anonymous):

if y varies inversely as x, find the inverse variation equation for the situations : y=5 when x=4 AND y=65 when x=9

OpenStudy (anonymous):

y = a/bx 5 = a/4b + c 20 = a/b + 4c 20b - 4cb = a 65 = (20b - 4cb)/9b + c 585 = 20-4c + 9c c = 113 20b - 4cb = a 20b - 452b = a -432b = a now you can choose whatever a and you will get the corresponding b.. y = a/bx + 113 , a = -432b